Pattison family considers the opportunity to finance the purchase of their first $500 thousand worth

single-family house in Santa Rosa, CA. To qualify for a 30-year fixed-rate mortgage, the down

payment should be 10% of the purchase price. To accumulate money for the down payment, Mr.

Pattison plans to invest $35,000 in the stock portfolio that earns 7.65% per annum. Assume that Mr.

Pattison reinvests annual gains at the same interest rate as the initial investment. How much money

will Mr. Pattison accumulate with this investment within five years? Is the initial investment

amount sufficient to accumulate in five years the required down payment under the above

mortgage? Round your answers to the nearest dollar.


1
Expert's answer
2021-11-17T11:04:33-0500

the down payment:

"500 000\times0.1=50 000"

dividends on shares:

"35 000\times0.0765=2677.5=2678"

if he puts in an annual dividend:

"FV=2677.5(1+0.0765)^5=3870.79=3871"

then he does not have enough money for the initial payment

"FV=35000(1+0.0765)^5=50598.57=50599"

if he invests 35,000,then he will get a down payment on the initial payment